OpenStudy (anonymous):

A Chemist has two saline solution. One is 2% salt, and the other is 6% salt. How much of each solution should be mixed to make 2 liters of saline solution that is 3% salt?

OpenStudy (anonymous):

Solution A has 0.02A salt and Solution B has 0.06B salt. If you mix A and B you will have: A+B=2 Liters and the amount of salt is: 0.02A+0.06B=0.03x2=0.06 Then you have two equations: A+B=2 and 0.02A+0.06B=0.06 Multiply the first one by 0.06 and you get: 0.06A+0.06B=0.12. Substract the second: (0.06-0.02)A+(0.06-0.06)B=0.12-0.06 Then 0.04A=0.06--->A=6/4=1.5 liters and B=2-1.5=0.5 liters Now check: 1.5 Liters of A and 0.5 Liters of B make two liters 1.5 liters of A mean 1.5 x 0.02=0.03 salt 0.5 liters of B mean 0.5 x 0.06 =0.03 salt Total salt=0.06 and concentration = 0.06/2=0.03 = 3%